Test the following numbers for divisibility by 6,9 and 11 ( do not divide or factories)



a. 6798340



b. 54786978

1
Expert's answer
2022-03-04T05:14:39-0500

a. 6798340

Divisibility by 6:

The number should be divisible by 2 and 3 both.

Num = 679840 is even number and therefore divisible by 2.

Sum = 6+7+9+8+4+0 = 34 is not divisible by 3, therefore, the number 6798340 is not divisible by 3.

Therefore, the number 6798340 is NOT DIVISIBLE by 6.

 

Divisibility by 9:

Sum = 6+7+9+8+4+0 = 34 is not divisible by 9, therefore, the number 6798340 is not divisible by 9.

 

Divisibility by 11:

Sum = 6-7+9-8+4-0 = 4 is not divisible by 11.

Hence the number is not divisible by 11.

 

b. 54786978

Divisibility by 6:

The number should be divisible by 2 and 3 both.

Num = 54786978 is even number and therefore divisible by 2.

Sum = 5+4+7+8+6+9+7+8 = 54 is divisible by 3, therefore, the number 54786978 is divisible by 3.

Therefore, the number 54786978 is DIVISIBLE by 6.

 

Divisibility by 9:

Sum = 5+4+7+8+6+9+7+8 = 54 is divisible by 9, therefore, the number 54786978 is divisible by 9.

 

Divisibility by 11:

Sum = 5-4+7-8+6-9+7-8=  -4 is not divisible by 11.

Hence the number is not divisible by 11.